• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Roedelbronn

    Log entry time 08:05:54 on June18,2002

    Entry number 86350

    keyword=Owl shift summary

    Crew: M. Roedelbronn, B. Reitz, H. Ibrahim

    Came in partway through kinematics Q2_f301

    00:28 Reset slow controls for HRS-R. When the crate came back up the VDC HV had both tripped. We tried to bring them back up to 4.0 kV but only the R_TOP was working. We tried to reset the slow controls once more and encountered the same problem.

    00:40 Took the Hall to Controled Access and entered the detector hut to try a local reset by cycling power to the power supply crate - also unsuccessful.

    01:30 Jack Segal came in and reset the power supply locally. He said our cycling of crate power should have had the same effect, and it wasn't clear why that didn't work. Jack's reset was successful, and the VDC came back up to 4 kV. He may look more closely at it during tomorrow's maintenance. He has experienced quirky behavior in resetting these types of power supplies before.

    02:00 Finished in the Hall, status switched to Beam Permit. Awaiting beam.

    02:15 We have beam back and have resumed taking data. The plan is to continue until 04:00 and then switch to the re-measurement of various Q2D_f00 points per log entry 86291.

    03:20 ROC14 crashed. Reset crate from Hall A Slow Controls

    04:00 Stopped data taking to move to Optics run (run 1612) and then to remeasurement of f00 points.

    04:57 Started the remeasurements a bit later than planned because it took a long time to change the fields of right arm Q2 and Q3. I had to manually set the values. Run 1613 is the optics run for the re-measurement.

    05:10 Switched to 15cm LH2 target, HRS-R sieve slit inserted. Runs 1614 and 1615 run with this configuration for a total time of ~60 minutes

    05:57 Beam down for q.e. measurement

    06:04 q.e. measurement complete. Beam coming back.

    06:19 Set HRS-R collimator to OPEN for a ~15 min run (1616 &1617 )

    06:45 Switching to 15cm LD2 target

    07:02 HRS-L Q1 tripped. Waiting for current to fall to zero...

    07:21 Q1 back. Starting run 1618.

    07:36 Starting run 1619, continuation of LD2 measurements. Since there is not really enough time to switch to the 4cm LH2 tgt, the plan is to keep with this configuration until beam goes away for the machine development day.

    07:48 Starting run 1620, continuation of LD2 measurements.

    08:01 Beam gone for machine development.